OpenStudy (anonymous):

help please 24z^5-12z^4 divided by 3z^2

OpenStudy (anonymous):



OpenStudy (akashdeepdeb):

\[ \frac{ 24z^5-12z^4 }{3z^2 } = \frac{24z^5 }{3z^2 }-\frac{ 12z^4 }{ 3z^2 } = 8z^3-4z^2\] Understood? :)

OpenStudy (anonymous):

yes, I had that answer but I didn't know if I was right or not

OpenStudy (akashdeepdeb):

:)